OpenStudy (anonymous):

The average of 11 numbers is 21 . The average of seven of those numbers is 25. What is the average of the remaining four numbers?

OpenStudy (anonymous):

Has anyone on here taken there SAT yet? I am taking a SAT prep test I found online and am a little concerned but am sure I will do well. Any advice for preparing?

OpenStudy (anonymous):



OpenStudy (anonymous):

SAT prep is more about understanding what they are askiing for....repetition is the only advice I can give you. My son just went through this last year.

OpenStudy (anonymous):

for this problem; y/7 = 25 y/4 = x so, (25 + x)/2 = 21 x =17 so the average of the other 4 number is 17

OpenStudy (anonymous):

how did you get the answer to that question?

OpenStudy (anonymous):

oh wait.. I see u already wrote it. :)

OpenStudy (anonymous):